### v1

# Upper bounds on the genus of Albanese fibrations

Abstract.
Let $S$ be a minimal irregular surface of general type, whose Albanese map induces a hyperelliptic fibration $f:\,S\to B$ of genus $g$. We prove a quadratic upper bound on the genus $g$, i.e., $g\leq h\big{(}\chi(\mathcal{O}_{S})\big{)}$, where $h$ is a quadratic function. We also construct examples showing that the quadratic upper bounds can not be improved to the linear ones. In the special case when $p_{g}(S)=q(S)=1$, we show that $g\leq 14$.
Key words and phrases:
surface of general type, Albanese map, fibration
2010 Mathematics Subject Classification:
14J29; 14J10; 14D06
This work is supported by National Natural Science Foundation of China, Fundamental Research Funds for central Universities, and Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014)
1. Introduction
We work over the complex number throughout this paper. Let $S$ be a minimal irregular surface of general type, and $a:\,S\to\mathrm{Alb}(S)$ its Albanese map. We are interested in the case when the image $a(S)$ is a curve. In this case, the Albanese map induces a fibration, which we call the Albanese fibration of $S$:
$f:\,S\longrightarrow B.$
In fact, by the universal property of the Albanese map, $B\cong a(S)$, and under this isomorphism the above fibration $f$ is nothing but the Albanese map of $S$. Let $g$ be the genus of a general fiber of $f$. A natural problem is to study the behavior of the genus $g$.
[…]
Can we give an upper bound on the genus $g$ of the Albanese fibration of $S$?
By [Cat00], it is known that the genus $g$ of the Albanese fibration of $S$ is a differential invariant, and hence is also a deformation invariant. Fixing $\chi(\mathcal{O}_{S})$ and $K_{S}^{2}$, there are only finitely many deformation equivalence classes of such surfaces, cf. [BHPV04, § VII]. Hence there must be an upper bound on $g$ depending on $\chi(\mathcal{O}_{S})$ or $K_{S}^{2}$. However, explicit upper bounds are only known under some extra assumptions.
  1. (1)
    Suppose that $g(B)>1$. Then $g\leq\frac{\chi(\mathcal{O}_{S})}{g(B)-1}+1\leq\chi({\mathcal{O}_{S}})+1$ by the semi-positivity of the Hodge bundle $f_{*}\mathcal{O}_{S}\big{(}K_{S/B}\big{)}$. Moreover, this bound is sharp since there are generalised hyperelliptic surfaces, whose Albanese map is a fibration of genus $g=\chi({\mathcal{O}_{S}})+1$ (cf. [Cat00]).
  2. (2)
    Suppose that $g(B)=1$ and $K_{S}^{2}<4\chi(\mathcal{O}_{S})$. From the slope inequality [CH88, Xia87a] it follows that $\frac{K_{S}^{2}}{\chi(\mathcal{O}_{S})}\geq\frac{4(g-1)}{g}$, and hence $g\leq\frac{4\chi(\mathcal{O}_{S})}{4\chi(\mathcal{O}_{S})-K_{S}^{2}}\leq 4\chi(\mathcal{O}_{S})$. Konno [Kon96] showed that $g\leq 6$ if moreover $S$ is an even canonical surface.
  3. (3)
    Suppose that $g(B)=1$ and $K_{S}^{2}=4\chi(\mathcal{O}_{S})$. Ishida [Ish05] proved that $g\leq 2\chi(\mathcal{O}_{S})+2$ if moreover $f$ is hyperelliptic.
All the upper bounds above are linear in $\chi(\mathcal{O}_{S})$. Our first aim is to give an upper bound on the genus $g$ of hyperelliptic Albanese fibrations of surfaces of general type with $K_{S}^{2}>4\chi(\mathcal{O}_{S})$.
Theorem 1.1.
Let $S$ be a minimal irregular surface of general type with $K_{S}^{2}>4\chi(\mathcal{O}_{S})$, and $f:\,S\to B$ its Albanese fibration whose general fiber is of genus $g$. Suppose that $f$ is hyperelliptic. Then
$g\leq h\big{(}\chi(\mathcal{O}_{S})\big{)},\qquad\text{where $h$ is some quadratic function.}$.
The above quadratic function can be explicitly written out, see Theorem 4.1. We will also construct in Section 5 examples of minimal irregular surfaces of general type with $K_{S}^{2}>4\chi(\mathcal{O}_{S})$ admitting hyperelliptic Albanese fibrations, whose general fiber is of genus $g$ as large as a quadratic function in $\chi(\mathcal{O}_{S})$. This is different from the phenomenons when $K_{S}^{2}\leq 4\chi(\mathcal{O}_{S})$, where the known upper bounds are all linear in $\chi(\mathcal{O}_{S})$ as mentioned above.
[…]
As mentioned above, the genus $g$ is a deformation invariant when $f$ is the Albanese fibration. Hence one may expect an upper bound on $g$ using invariants of $S$. In general, there is no such upper bound on $g$ if $f$ is not the Albanese fibration. In fact, Penegini-Polizzi ([PP17]) constructed a minimal surface $S$ with $p_{g}(S)=q(S)=2,K_{S}^{2}=4$ (whose Albanese map is generically finite), on which there are fibrations $f_{k}:\,S\to B$, such that the genera of $f_{k}$’s can be arbitrarily large. Nevertheless, in the case when $f:\,S\to B$ is hyperelliptic, the quadratic upper bound proved in Theorem 1.1 holds true even if $f$ is not the Albanese fibration, and hence our results can be applied in a more general situation. See Theorem 4.1 for more details.
Our another interest is the classification of minimal irregular surface of general type with $\chi(\mathcal{O}_{S})=1$. Although the value $\chi(\mathcal{O}_{S})$ is the minimal possibility among surfaces of general type, the classification of such surfaces is widely open. Beauville proved in an appendix to [Deb82] that $p_{g}(S)=q(S)\leq 4$ for such surfaces, and the equality holds if and only if $S$ is isomorphic to a product of two curves of genus $2$. We are more interested in the case when the Albanese map of $S$ induces a fibration. In this case, $p_{g}(S)=q(S)\leq 3$ by Beauville’s result above. In a series of works [HP02, Pen11, Pir02, Zuc03], a full classification has been obtained when $p_{g}(S)=q(S)=3$ or $p_{g}(S)=q(S)=2$. However, the case when $p_{g}(S)=q(S)=1$ seems to be much more mysterious. An explicit upper bound on the genus $g$ of the Albanese fibration is still unknown, although theoretically such an upper bound exists since there is at most finitely many deformation equivalence classes of such surfaces. We prove an explicit upper bound on the genus if the Albanese fibration is hyperelliptic.
Theorem 1.2.
Let $S$ be a minimal irregular surface of general type with $p_{g}(S)=q(S)=1$. Suppose that the Albanese fibration $f:\,S\to B$ is hyperelliptic of genus $g$. Then $g\leq 14$. More precisely, it holds $g\leq 10$ except for the following two possible cases:     (i) $K^{2}=8,~{}g=14$;     (ii) $K^{2}=8,~{}g=11$.
The known examples of surfaces with $p_{g}(S)=q(S)=1$ are mostly restricted to the region $g\leq K_{S}^{2}$. The first example with $g>K_{S}^{2}$ was constructed in [FP15], in which $g=7$ and $K_{S}^{2}=6$. More recently, An example with $g=19$ and $K_{S}^{2}=9$ was constructed in [CKY17]. By our result above, the Albanese fibration of the example in [CKY17] must be non-hyperelliptic. The paper is organized as follows. In Section 2, we mainly review some basic facts about surface fibrations and do some technical preparations. In Section 3 and Section 4, we prove Theorem 1.2 and Theorem 1.1 respectively. Finally in Section 5, we construct examples of irregular surfaces of general type with hyperelliptic fibrations of genus $g$ as large as a quadratic function in $\chi(\mathcal{O}_{S})$.
2. Preliminaries
In this section, we mainly review some basic facts and fix the notations. In Section 2.1, we recall some general facts about the surface fibrations, and refer to [BHPV04] for more details; and then in Section 2.2 we restrict ourselves to the theory on the hyperelliptic fibrations, which are mainly due to Xiao [Xia91, Xia92].
2.1. The surface fibrations
Let $f:S\rightarrow B$ be a fibration of curves of genus $g\geq 2$, i.e. $f$ is a proper morphism from a smooth projective surface $S$ onto a smooth projective curve $B$ with connected fibers over the complex number, and the general fiber is a smooth projective curve of genus $g$. The fibration $f$ is called relatively minimal if there is no $(-1)$-curve (i.e. a smooth rational curve with self-intersection $-1$) contained in the fiber of $f$. It is called hyperelliptic if its general fiber is hyperelliptic, smooth if all its fibers are smooth, isotrivial if all its smooth fibers are isomorphic to each other, and locally trivial if it is both smooth and isotrivial. Let $\omega_{S}$ (resp. $K_{S}$) be the canonical sheaf (resp. canonical divisor) of $S$, and let $\omega_{S/B}=\omega_{S}\otimes f^{*}\omega_{B}^{\vee}$ (resp. $K_{f}=K_{S/B}=K_{S}-f^{*}K_{B}$) the relative canonical sheaf (resp. the relative canonical divisor) of $f$. Put $b:=g(B)$, $p_{g}:=h^{0}(S,\omega_{S})$, $q:=h^{1}(S,\omega_{S})$, $\chi=\chi(\mathcal{O}_{S}):=p_{g}-q+1$, and let $\chi_{top}(S)$ be the topological Euler characteristic of $S$. The basic invariants of $f$ are:
$\chi_{f}$ $\,=\chi-(g-1)(b-1);$
$K_{f}^{2}$ $\,=K_{S}^{2}-8(g-1)(b-1);$
$e_{f}$ $\,=\chi_{top}(S)-4(g-1)(b-1).$
These invariants satisfy the following properties:
  1. (1)
    $\chi_{f}=\deg f_{*}\omega_{S/B}$ is the degree of the Hodge bundle $f_{*}\omega_{S/B}$. Moreover $\chi_{f}\geq 0$, and the equality holds if and only if $f$ is locally trivial.
  2. (2)
    When $f$ is relatively minimal, $K_{f}^{2}\geq 0$, and the equality holds if and only if $f$ is locally trivial.
  3. (3)
    $e_{f}=\sum e_{F}$, where $e_{F}:=\chi_{\rm top}(F_{red})-(2-2g)$ for any fiber $F$ and $F_{red}$ is the reduced part of $F$. Moreover, $e_{F}\geq 0$, and the equality holds if and only if $F$ is smooth. Hence $e_{f}\geq 0$, and $e_{f}=0$ iff $f$ is smooth.
  4. (4)
    The above three invariants satisfy the Noether equality: $12\chi_{f}=K_{f}^{2}+e_{f}$.
Suppose that $f$ is relatively minimal and not locally trivial. Then both $K_{f}^{2}$ and $\chi_{f}$ are strictly positive. In this case, the slope of $f$ is defined to be $\lambda_{f}=\frac{K_{f}^{2}}{\chi_{f}}$. According to the non-negativity of these basic invariants of $f$, it holds $0<\lambda_{f}\leq 12$. The so-called slope inequality, proved independently by Cornalba-Harris [CH88] and Xiao [Xia87a], states that
$\lambda_{f}\geq\frac{4(g-1)}{g}.$ (2.1)
2.2. Invariants of hyperelliptic fibrations
In this subsection, we restrict to the hyperelliptic fibrations. Let $f:S\rightarrow B$ be a relatively minimal hyperelliptic fibration. The relative canonical map of $f$ is generically of degree 2. Hence it determines an involution $\sigma$ on $S$ whose restriction on a general fiber $F$ of $f$ is the hyperelliptic involution of $F$. Let $\mu:\tilde{S}\rightarrow S$ be the blowing-ups of all the isolated fixed points of $\sigma$ and let $\tilde{\sigma}$ be the induced involution on $\tilde{S}$. The quotient space $\tilde{P}=\tilde{S}/\langle\tilde{\sigma}\rangle$ is a smooth surface, and $f$ induces a ruling $\tilde{h}:\tilde{P}\rightarrow B$ on $\tilde{P}$. Also, the quotient map $\tilde{\pi}:\tilde{S}\rightarrow\tilde{P}$ is a double cover which is determined by the pair $(\tilde{R},\tilde{\delta})$, where $\tilde{R}$ is the branch locus of $\tilde{\pi}$ and $\tilde{\delta}$ is a line bundle such that $\mathcal{O}_{\tilde{P}}(\tilde{R})\cong\tilde{\delta}^{\otimes 2}$ and $\tilde{\pi}_{*}\mathcal{O}_{\tilde{S}}\cong\mathcal{O}_{\tilde{P}}\oplus\tilde{\delta}^{\vee}$. For any contraction $\varphi:\tilde{P}\rightarrow P^{\prime}$ and $R^{\prime}=\varphi(\tilde{R})$, the double cover $\tilde{\pi}$ induces a double cover $S^{\prime}\rightarrow P^{\prime}$, which is determined by $(R^{\prime},\delta^{\prime})$. We call $(R^{\prime},\delta^{\prime})$ the image of $(\tilde{R},\tilde{\delta})$.
[…]
There exists a contraction of a ruled surfaces $\psi:\tilde{P}\rightarrow P$:
[…]
such that $P$ is a geometrical ruled surface (i.e. any fiber of $h$ is $\mathbb{P}^{1}$), the singularities of $R$ are at most of multiplicity $g+2$, and the self-intersection $R^{2}$ is the smallest among all such choices, where $(R,\delta)$ is the image of $(\tilde{R},\tilde{\delta})$ in $P$.
We remark that the minimality of $R^{2}$ implies that $s_{g+2}=0$ if $g$ is even (c.f. [LZ17, Thm 2.6 and its proof]) The contraction $\psi$ in Section 2.2 can be decomposed into $\tilde{\psi}:\tilde{P}\rightarrow\hat{P}$ and $\hat{\psi}:\hat{P}\rightarrow P$, where $\tilde{\psi}$ and $\hat{\psi}$ are composed respectively of resolutions of negligible and non-negligible singularities of $R$, cf. [Xia91, Def 4] or [LZ17, Def 2.1]. Let $(\hat{R},\hat{L})$ be the image of $(\tilde{R},\tilde{L})$ in $\hat{P}$. Let $\hat{\psi}=\hat{\psi}_{1}\circ\dotsm\circ\hat{\psi}_{t}$ be the decomposition of $\hat{\psi}$, where $\hat{\psi}_{i}:\hat{P}_{i}\rightarrow\hat{P}_{i-1}$ is a blowing-up at $y_{i-1}$, $\hat{P}_{0}=P$ and $\hat{P}_{t}=\hat{P}$. Let $\hat{R}_{i}$ be the image of $\hat{R}$ in $\hat{P}_{i}$. We remark that the decomposition of $\hat{\psi}$ is not unique. If $y_{i-1}$ is the singular point if $\hat{R}_{i-1}$ of odd multiplicity $2k+1~{}(k\geq 1)$ and there is a unique singular point $y$ of $\hat{R}_{i}$ on the exceptional curve ^$\hat{\mathcal{E}}_{i}$ of multiplicity $22k+$ 2, then we always assume that $\hat{\psi}_{i+1}:\hat{P}_{i+1}\rightarrow\hat{P}_{i}$ is the standard blowing-up at $y_{i}=y$. We call such a pair $(y_{i-1},y_{i})$ a singularity of type $(2k+1\rightarrow 2k+1)$ and call $y_{i-1}$ (resp. $y_{i}$) the first (resp. second) component of such a singularity.
[…]
For any singular fiber $F$ of $f$ and $23\leq i\leq g+$ 2, the $i$-th singularity index of $F$ is defined as follows (with respect to the contraction $\psi$): (1) if $i$ is odd, $s_{i}(F)$ equals the number of $(i\rightarrow i)$ type singularities of $R$ over the image $f(F)$; (2) if $i$ is even, $s_{i}(F)$ equals the number of singularities of multiplicity $i$ or $i+1$ of $R$ over the image $f(F)$, neither belonging to the second component of $(i-1\rightarrow i-1)$ type singularities nor to the first component of $(i+1\rightarrow i+1)$ type singularities.
Let $K_{\hat{P}/B}=K_{\hat{P}}-h^{*}K_{B}$ and $R^{\prime}={\hat{R}}\setminus{\hat{V}}$, where $\hat{V}$ is the union of isolated vertical $(-2)$-curves in $\hat{R}$. Here a curve $C\subset\hat{R}$ is said to be isolated in $R$ if there is no other curve $C^{\prime}\subset\hat{R}$ such that $C\cap C^{\prime}\neq\emptyset$. We define
$s_{2}\triangleq(K_{\hat{P}/B}+R^{\prime})\cdot R^{\prime}\qquad\text{and}\qquad s_{i}\triangleq\sum_{F~{}\text{is singular}}s_{i}(F),~{}~{}3\leq i\leq g+2.$
Note that $s_{i}$ is nonnegative for $i\geq 3$ by definition, but it is not clear whether $s_{2}$ is nonnegative or not.
[…]
These singularity indices $s_{i}$’s defined above are independent on the choices of $\psi$ in Section 2.2.
On the other hand we can write $R\sim-(g+1)K_{P/B}+nF$, where $\sim$’ stands for numerical equivalence, and $n=\frac{R^{2}}{4(g+1)}$ is an integer. The following formulas for hyperelliptic fibrations are due to Xiao; we refer to [Xia91, Thm 1] or [LZ17, Thm 2.6] for a proof.
Theorem 2.1.
Let $f:S\rightarrow B$ be a hyperelliptic fibration of genus $g\geq 2$, and let $s_{i}$’s be the singularity indices as above. Then
$\chi_{f}$ $\,=\frac{1}{2}gn-\sum_{k=1}^{[\frac{g+1}{2}]}k^{2}s_{2k+1}-\sum_{k=2}^{[\frac{g+1}{2}]}\frac{k(k-1)}{2}s_{2k}$
$\,=\frac{gs_{2}+(g^{2}-2g-1)s_{g+2}}{4(2g+1)}+\sum_{k=1}^{[\frac{g}{2}]}\frac{k(g-k)}{2g+1}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{k(g-k+1)}{2(2g+1)}s_{2k};$
$K_{f}^{2}$ $\,=(2g-2)n+s_{g+2}-\sum_{k=1}^{[\frac{g+1}{2}]}(2k-1)^{2}s_{2k+1}-\sum_{k=2}^{[\frac{g+1}{2}]}2(k-1)^{2}s_{2k}$
$\,=\frac{(g-1)\big{(}s_{2}+(3g+1)s_{g+2}\big{)}}{2g+1}+\sum_{k=1}^{[\frac{g}{2}]}\frac{12k(g-k)-2g-1}{2g+1}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{6k(g-k+1)-4g-2}{2g+1}s_{2k};$
$e_{f}$ $\,=s_{2}-2s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}s_{2k+1}+2\sum_{k=2}^{[\frac{g+1}{2}]}s_{2k}.$
Suppose that $f$ is locally non-trivial. Then the above formulas implies in particular that
$\left(\lambda_{f}-\frac{4(g-1)}{g}\right)\chi_{f}=K_{f}^{2}-\frac{4(g-1)}{g}\chi_{f}$ (2.2)
$=\frac{g^{2}-1}{g}s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}\frac{4k(g-k)-g}{g}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{2k(g-k+1)-2g}{g}s_{2k}.$
[…]
Let $f:S\rightarrow B$ be a locally non-trivial hyperelliptic fibration of genus $g\geq 2$. Then the slope $\lambda_{f}$ satisfies
$\frac{4(g-1)}{g}\leq\lambda_{f}\leq\left\{\begin{aligned} &12-\frac{8g+4}{g^{2}},&~{}&\text{~{}if $g$ is even}\\ &12-\frac{8g+4}{g^{2}-1},&&\text{~{}if $g$ is odd}\end{aligned}\right\}<12.$ (2.3)
If moreover the base curve $B$ is an elliptic curve, then the upper bound of $\lambda_{f}$ can be improved. In this case, $g(B)=1$ and hence the relative invariants of $f$ equal the corresponding invariants of the surface $S$, i.e.,
$K_{f}^{2}=K_{S}^{2},\qquad\chi_{f}=\chi=\chi(\mathcal{O}_{S}),\qquad e_{f}=\chi_{top}(S).$ (2.4)
By the Miyaoka-Yau inequality together with (2.4), one obtains
$\lambda_{f}=\frac{K_{f}^{2}}{\chi_{f}}\leq 9.$ (2.5)
To end this section, we mention that by (2.2)together with (2.5) and (2.3), one obtains
$\chi_{f}$$\geq$ $\frac{g^{2}-1}{5g+4}s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}\frac{(4k-1)g-4k^{2}}{5g+4}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{2(k-1)(g-k)}{5g+4}s_{2k},\quad\text{if~{}}g(B)=1,$ (2.6)
$\chi_{f}$ $>$ $\frac{g^{2}-1}{8g+4}s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}\frac{(4k-1)g-4k^{2}}{8g+4}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{2(k-1)(g-k)}{8g+4}s_{2k},\quad\text{if~{}}g(B)\geq 2.\qquad\quad$ (2.7)
3. Upper bound on hyperelliptic Albanese fibrations with $p_{g}=q=1$
The aim in this section is to prove Theorem 1.2. So we always assume in this section that $S$ is a minimal irregular surface of general type with $p_{g}=q=1$ and its Albanese fibration $f:S\rightarrow B$ is hyperelliptic. In this case, $g(B)=q=1$ and $f$ must be locally non-trivial. By (2.4), the relative invariants equal the corresponding invariants of $S$. We use singularity indices $s_{i}$’s introduced in Section 2.2. Since it is unknown whether $s_{2}$ is nonnegative, we divide the proof into two cases, depending on whether $s_{2}<0$ or $s_{2}\geq 0$. The proof of Theorem 1.2 will be completed in Section 3, Section 3 and Section 3.
[…]
Let $f:\,S\to B$ be as in Theorem 1.2. If $s_{2}<0$, then $g\leq 5$.
Proof.
We first claim that
[…]
If $s_{2}<0$, then we have $K_{S}^{2}\leq 7$.
Proof of Section 3.
If $s_{2}<0$, using notations in Section 2.2, we see that $\hat{R}$ must contain some isolated curve $C$ with $C^{2}\neq-2$. So there must be some smooth rational curves $\widetilde{C}$ contained in fibers of $f$. Assume $\widetilde{C}^{2}=-n$, by [Miy84, Theorem 1.1] we have
$\frac{3}{2}\leq\frac{(n+1)^{2}}{3n}\leq\chi_{\rm top}(S)-\frac{1}{3}K_{S}^{2}=\frac{4}{3}(9\chi-K_{S}^{2}).$
Hence we get $K_{S}^{2}\leq 9-\frac{9}{8}$, i.e. $K_{S}^{2}\leq 7$ since $K_{S}^{2}$ is an integer. ∎
Come back to the proof of Section 3. By Theorem 2.1, we have
$K_{f}^{2}-\frac{8g-14}{g-1}\chi_{f}$
=\,= $-\frac{g-2}{2(g-1)}s_{2}+\frac{g^{2}+2g-5}{2(g-1)}s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}\left(\frac{2k(g-k)}{g-1}-1\right)s_{2k+1}+\sum_{k=3}^{[\frac{g+1}{2}]}\left(\frac{k(g-k+1)}{g-1}-2\right)s_{2k}$
[…]
Hence
$\frac{g-2}{2(g-1)}\leq K_{f}^{2}-\frac{8g-14}{g-1}\chi_{f}\leq 7-\frac{8g-14}{g-1}=\frac{7-g}{g-1}.$
It follows that $g\leq\frac{16}{3}$, i.e., $g\leq 5$ as required. ∎
In the rest part of this section, we assume that $s_{2}\geq 0$. We first claim that
[…]
Let $f:\,S\to B$ be as in Theorem 1.2. Suppose that $s_{2}\geq 0$. Then
$&\text{if $g\geq 6$,}&\quad&\text{then $s_{g+2}=0$;}\\ &\text{if $g\geq 11$,}&&\text{then $s_{2k+1}=0,\quad\forall~{}k\geq 2$;}\\ &\text{if $g\geq 13$,}&&\text{then $s_{2k}=0,\quad\forall~{}k\geq 6$.}\end{aligned}\right.$ (3.1)
Proof.
If $s_{g+2}>0$, then by (2.4) and (2.6) we have
$1=\chi=\chi_{f}\geq\frac{g^{2}-1}{5g+4}s_{g+2}\geq\frac{g^{2}-1}{5g+4},\quad\Longrightarrow\quad g^{2}-5g-5\leq 0$
Hence $g<6$. This shows that $s_{g+2}=0$ if $g\geq 6$. The rest two inequalities can be proved similarly using (2.6), and left to the readers. ∎
[…]
Let $f:\,S\to B$ be as in Theorem 1.2. If $g\geq 13$, then there is only one possible case for $(K_{S}^{2},g)$: $K_{S}^{2}=8,g=14$. In particular, $g\leq 14$.
Proof.
Since $g\geq 13$, by (3.1) together with Theorem 2.1,
$1\displaystyle 2g+$ 1 $=$ $\frac{g}{4}s_{2}+(g-1)s_{3}+(g-1)s_{4}+\frac{3}{2}(g-2)s_{6}+2(g-3)s_{8}+\frac{5}{2}(g-4)s_{10},$ (3.2)
$\frac{gn}{2}-1$ $=$ $s_{3}+s_{4}+3s_{6}+6s_{8}+10s_{10}.$ (3.3)
(1) We show first that $s_{10}=0$. Otherwise by (3.2), $s_{10}=1$ and
$11-\frac{g}{2}=\frac{g}{4}s_{2}+(g-1)s_{3}+(g-1)s_{4}+\frac{3}{2}(g-2)s_{6}+2(g-3)s_{8}.$
Since $g\geq 13$ and each $s_{i}$ is a nonnegative integer, it follows that $s_{3}=s_{4}=s_{6}=s_{8}=0$, and hence $11-\frac{g}{2}=\frac{g}{4}s_{2}$, i.e., $s_{2}+2=\frac{44}{g}$, which implies that $s_{2}=0$ and $g=22$. Substitute into the formulas of Theorem 2.1, one obtains that $K_{S}^{2}=K_{f}^{2}=10>9=9\chi(\mathcal{O}_{S})$, which is a contradiction. (2) We show that if $s_{8}\neq 0$, then
$s_{8}=1,~{}s_{6}=s_{4}=s_{3}=0,~{}s_{2}=2,~{}g=14,K_{S}^{2}=8.$
In fact, By (1), we already prove $s_{10}=0$. According to (3.2), $s_{8}=1$ and
$7=\frac{g}{4}s_{2}+(g-1)s_{3}+(g-1)s_{4}+\frac{3}{2}(g-2)s_{6}.$
Since $g\geq 13$, it follows that $s_{3}=s_{4}=s_{6}=0$, and $gs_{2}=28$. Hence either $g=28,s_{2}=1$ or $g=14,s_{2}=2$. The first case is impossible; otherwise, by (3.3) one has $n=\frac{1}{2}$, which is a contradiction since $n$ is an integer. Finally, $K_{S}^{2}=8$ follows from the formula in Theorem 2.1. (3) We show that $s_{6}=0$. Otherwise we would have $s_{6}=1$ again by (3.2). Moreover, by the arguments (1) and (2) above, we have $s_{10}=s_{8}=0$. Moreover, we have
$\frac{g}{2}+4=(2g+1)-\frac{3}{2}(g-2)=\frac{g}{4}s_{2}+(g-1)s_{3}+(g-1)s_{4}.$
Hence $s_{3}=s_{4}=0$, $s_{2}=3$ and $g=16$. Substitute into (3.3), one obtains that $n=\frac{1}{2}$, which gives a contradiction as $n$ is an integer. (4) We show that $s_{3}=s_{4}=0$. Otherwise by (2), $s_{8}=0$. Combining with (1) and (2) above, we have $s_{6}=s_{8}=s_{10}=0$, and hence by (3.2)
$2g+1=\frac{g}{4}s_{2}+(g-1)(s_{3}+s_{4}).$
Hence $s_{3}+s_{4}\leq 2$. If $s_{3}+s_{4}=2$, then $gs_{2}=12$, which is impossible since $g\geq 13$. If $s_{3}+s_{4}=1$, then by (3.3) one gets $gn=4$, which is impossible since $g\geq 13$ and $n\geq 1$. In conclusion, if $g\geq 13$, then except the possible case $(K_{S}^{2},g)=(8,14)$, one has $s_{i}=0$ for all $i>2$. Hence by (3.3) it holds $gn=2$, which is impossible since $n$ is an integer. This completes the proof. ∎
[…]
Let $f:\,S\to B$ be as in Theorem 1.2.
  1. (1)
    The genus $g\neq 12$.
  2. (2)
    If $g=11$, then $K_{S}^{2}=8$.
Proof.
(1) Suppose that $g=12$. Then by (3.1) together with Theorem 2.1,
$25=2g+1=3s_{2}+11s_{3}+11s_{4}+15s_{6}+18s_{8}+20s_{10}+21s_{12}.$
Note that all the $s_{i}$’s are non-negative integers. It is not difficult to show that $s_{6}=s_{8}=s_{10}=s_{12}=0$. Thus $253s_{2}+11(s_{3}+s_{4})=$ 25 . It follows that $s_{2}=1$ and $s_{3}+s_{4}=2$. By Theorem 2.1 one has $36n=\frac{gn}{2}=1+s_{3}+s_{4}=$ 3, which is impossible since $n$ is an integer. (2) Suppose that $g=11$. Similar as above, by (3.1) together with Theorem 2.1,
$46=2(2g+1)=\frac{11}{2}s_{2}+20s_{3}+20s_{4}+27s_{6}+32s_{8}+35s_{10}+36s_{12}.$
In particular, $s_{2}$ is even. If $s_{2}\geq 6$, then
$20s_{3}+20s_{4}+27s_{6}+32s_{8}+35s_{10}+36s_{12}\leq 13$
This is impossible since all $s_{i}$’s are non-negative integers. If $s_{2}=4$, then
$2420s_{3}+20s_{4}+27s_{6}+32s_{8}+35s_{10}+36s_{12}=$ 24 .
It is again impossible. If $s_{2}=2$, then
$3520s_{3}+20s_{4}+27s_{6}+32s_{8}+35s_{10}+36s_{12}=$ 35 .
Hence $s_{10}=1$ and $s_{i}=0$ for other $i>2$. Combining with Theorem 2.1, one computes $K_{S}^{2}=8$. If $s_{2}=0$, then
$4620s_{3}+20s_{4}+27s_{6}+32s_{8}+35s_{10}+36s_{12}=$ 46 .
One checks again this is impossible since all $s_{i}$’s are non-negative integers. This completes the proof. ∎
[…]
Let $f:\,S\to B$ be as in Theorem 1.2. By a similar argument as above, one can show that $g\neq 9$. More precisely, the list of all possibilities for $(K_{S}^{2},g)$ is as follow:
${}K_{S}^{2}~{}$ $g$
$9$ $4,~{}6,~{}8,~{}10$
$8$ $3,~{}4,~{}5,~{}6,~{}7,~{}8,~{}10,~{}11,~{}14~{}$
$7$ $3,~{}4,~{}5,~{}6$
$6$ $2,~{}3,~{}4,~{}5,~{}6,~{}7,~{}8$
$5$ $2,~{}3,~{}4$
$4$ $2,~{}3,~{}4$
${}3,~{}2~{}$ $2$
If $K_{S}^{2}=2$, then $g=2$ by the slope inequality (2.1). If $K_{S}^{2}=3$, then one can only get $g\leq 4$ by the slope inequality. Catanese-Ciliberto [CC91, Prop 5.6 & Thm 5.7] proved that in this case $g=2$. Our method here provides a different proof of this result. Finally, it would be interesting to construct examples of hyperelliptic Albanese fibrations (or exclude the existence) in the above list $(K_{S}^{2},g)$.
4. Quadratic upper bounds on hyperelliptic Albanese fibrations
The main purpose of this section is to prove Theorem 1.1. We will prove the following upper bounds (using the relative invariants) on the genus $g$ for any locally non-trivial hyperelliptic fibration, not necessarily the Albanese fibration. We remark once more that the relative invariants of $f$ equal the corresponding invariants of the surface $S$ (cf. (2.4)) if the base $B$ is an elliptic curve.
Theorem 4.1.
Let $f:\,S\to B$ be a relatively minimal hyperelliptic fibration of genus $g\geq 2$. Suppose that $f$ is locally non-trivial.
  1. (1)
    If $\lambda_{f}<4$, then $g\leq\frac{4\chi_{f}}{4\chi_{f}-K_{f}^{2}}$.
  2. (2)
    If $\lambda_{f}=4$, then $g\leq 2\chi_{f}+2$.
  3. (3)
    If $\lambda_{f}>4$, then
    $g\leq\left\{\begin{aligned} &2\chi_{f}+1,&&\text{if $g(B)=0$};\\ &\frac{25}{2}\chi_{f}^{2}+14\chi_{f}+4,&\quad&\text{if $g(B)=1$};\\ &32\chi_{f}^{2}+20\chi_{f}+3,&&\text{if $g(B)\geq 2$}.\end{aligned}\right.$
We first prove Theorem 1.1 based on Theorem 4.1.
Proof of Theorem 1.1.
Let $f:\,S\to B$ be the Albanese fibration of $S$ and $g$ the genus of a general fiber. Then $g(B)=q\geq 1$. If $g(B)\geq 2$, by the non-negativity of the Hodge bundle, one obtains
$\chi(\mathcal{O}_{S})-(g-1)(g(B)-1)=\chi_{f}\geq 0.$
Hence $g\leq\frac{\chi(\mathcal{O}_{S})}{g(B)-1}+1\leq\chi({\mathcal{O}_{S}})+1$. If $g(B)=1$, then $f$ must locally non-trivial; otherwise $\chi(\mathcal{O}_{S})=\chi_{f}=0$ which is absurd. Thus by (2.4) together with Theorem 4.1 (3), one obtains a quadratic upper bound on the genus $g$. ∎
We next come back to the proof of Theorem 4.1. The proof relays heavily on the formulas given by Xiao in Theorem 2.1.
Proof of Theorem 4.1.
The first statement follows from the slope inequality (2.1). The second one was proved by Ishida [Ish05]. Here we can give a short proof based on Theorem 2.1. In fact, since $\lambda_{f}=4$, i.e., $K_{f}^{2}=4\chi_{f}$, by Theorem 2.1 one gets
$\chi_{f}$ $\,=\frac{g}{4}\left(K_{f}^{2}-\frac{4(g-1)}{g}\chi_{f}\right)$
$\,=\frac{g^{2}-1}{4}s_{g+2}+\sum_{k=1}^{[\frac{g}{2}]}\frac{(4k-1)g-4k^{2}}{4}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{(k-1)(g-k)}{2}s_{2k}$
$\,\geq\min\left\{\frac{g^{2}-1}{4},~{}\min_{1\leq k\leq[\frac{g}{2}]}\frac{(4k-1)g-4k^{2}}{4},~{}\min_{2\leq k\leq[\frac{g+1}{2}]}\frac{(k-1)(g-k)}{2}\right\}$
$\,=\frac{g-2}{2}.$
Thus $g\leq 2\chi_{f}+2$ as required.
It remains to prove (3). If $g(B)=0$, then
$\chi_{f}=\chi(\mathcal{O}_{S})+(g-1)=p_{g}-q+g\geq\frac{g-1}{2}.$
The inequality above follows from the inequality $q\leq\frac{g+1}{2}$ by Xiao [Xia87b] and the non-negativity of $p_{g}$. Hence $g\leq 2\chi_{f}+1$ as required. We consider next the case when $g(B)=1$, and we are going to prove by contradiction. Suppose that
$g>\frac{25}{2}\chi_{f}^{2}+14\chi_{f}+4.$ (4.1)
Then it is clear that $s_{g+2}=0$ by (2.6). Let
$k_{0}=\max\left\{k~{}\big{|}~{}s_{2k}\neq 0,\text{~{}or~{}}s_{2k+1}\neq 0\right\}.$
Then by (2.6) one obtains that
$k_{0}\leq\frac{5\chi_{f}+2}{2}.$ (4.2)
In fact, suppose on the contrary. Then $k_{0}\geq\frac{5\chi_{f}+3}{2}$, and hence from (2.6) it follows that
$\chi_{f}\geq\min\left\{\frac{(4k_{0}-1)g-4k_{0}^{2}}{5g+4},~{}\frac{2(k_{0}-1)(g-k_{0})}{5g+4}\right\}\geq\frac{\big{(}5\chi_{f}+1)\cdot\left(g-\frac{5\chi_{f}+3}{2}\right)}{5g+4}.$
It gives a contradiction to (4.1). Let
$r=\frac{k_{0}(k_{0}-1)}{g},\quad t=\lambda_{f}-\frac{4(g-1)}{g},\quad a_{k}=\frac{4k(g-k)-g}{g},\quad b_{k}=\frac{2k(g-k+1)-2g}{g}.$
Then by (2.2),
$\big{(}t\chi_{f}\big{)}^{2}+(3+4r)t\chi_{f}+4r^{2}+6r$ (4.3)
$=$ $\left(\sum_{k=1}^{k_{0}}a_{k}s_{2k+1}+\sum_{k=2}^{k_{0}}b_{k}s_{2k}\right)^{2}+(3+4r)\left(\sum_{k=1}^{k_{0}}a_{k}s_{2k+1}+\sum_{k=2}^{k_{0}}b_{k}s_{2k}\right)+4r^{2}+6r$
$\geq$ $\left(\sum_{k=1}^{k_{0}}a_{k}^{2}s_{2k+1}^{2}+\sum_{k=2}^{k_{0}}b_{k}^{2}s_{2k}^{2}\right)+(3+4r)\left(\sum_{k=1}^{k_{0}}a_{k}s_{2k+1}+\sum_{k=2}^{k_{0}}b_{k}s_{2k}\right)+4r^{2}+6r$
$=$ $\sum_{k=1}^{k_{0}}\big{(}a_{k}^{2}s_{2k+1}+(3+4r)a_{k}\big{)}s_{2k+1}+\sum_{k=2}^{k_{0}}\big{(}b_{k}^{2}s_{2k}+(3+4r)b_{k}\big{)}s_{2k}+4r^{2}+6r$
$\geq$ $\sum_{k=1}^{k_{0}}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{k_{0}}(4k^{2}-2k-2)s_{2k}.$
The last inequality above holds because
{$&a_{k}^{2}s_{2k+1}+(3+4r)a_{k}\geq 8k^{2}+4k-1,&~{}&\text{if $s_{2k+1}\neq 0$ with $1\leq k\leq k_{0}$};\\ &b_{k}^{2}s_{2k}+(3+4r)b_{k}\geq 4k^{2}-2k-2,&&\text{if $s_{2k}\neq 0$ with $2\leq k\leq k_{0}-1$};\\ &\big{(}b_{k_{0}}^{2}s_{2k_{0}}+(3+4r)b_{k_{0}}\big{)}s_{2k_{0}}+4r^{2}+6r\geq(4k_{0}^{2}-2k_{0}-2)s_{2k_{0}},&&\text{if $s_{2k_{0}}\neq 0$}.\end{aligned}\right.$
On the other hand, by (4.2) together with Theorem 2.1,
$\chi_{f}$ $\,=\frac{1}{8}s_{2}+\sum_{k=1}^{[\frac{g}{2}]}\frac{k}{2}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{k}{4}s_{2k}-\frac{1}{2g+1}\left(\frac{1}{8}s_{2}+\sum_{k=1}^{[\frac{g}{2}]}\frac{2k^{2}+k}{2}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{2k^{2}-k}{4}s_{2k}\right)$
$\,=\frac{1}{8}s_{2}+\sum_{k=1}^{k_{0}}\frac{k}{2}s_{2k+1}+\sum_{k=2}^{k_{0}}\frac{k}{4}s_{2k}-\frac{1}{2g+1}\left(\frac{1}{8}s_{2}+\sum_{k=1}^{k_{0}}\frac{2k^{2}+k}{2}s_{2k+1}+\sum_{k=2}^{k_{0}}\frac{2k^{2}-k}{4}s_{2k}\right);$
$s_{2}$ $\,=e_{f}+2s_{g+2}-\sum_{k=1}^{[\frac{g}{2}]}s_{2k+1}-2\sum_{k=2}^{[\frac{g+1}{2}]}s_{2k}=e_{f}-\sum_{k=1}^{k_{0}}s_{2k+1}-2\sum_{k=2}^{k_{0}}s_{2k}.$
Hence
$e_{f}+\sum_{k=1}^{k_{0}}(4k-1)s_{2k+1}+\sum_{k=2}^{k_{0}}2(k-1)s_{2k}-8\chi_{f}$
=\,= $\frac{1}{2g+1}\bigg{(}e_{f}+\sum_{k=1}^{k_{0}}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{k_{0}}(4k^{2}-2k-2)s_{2k}\bigg{)}.$
Note that the left hand of the above equality is an integer, and that the right hand of the above equality is positive. Therefore,
$e_{f}+\sum_{k=1}^{k_{0}}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{k_{0}}(4k^{2}-2k-2)s_{2k}\geq 2g+1$
Combining with (4.3), one obtains
$\big{(}t\chi_{f}\big{)}^{2}+(3+4r)t\chi_{f}+4r^{2}+6r+(12-\lambda_{f})\chi_{f}$
=\,= $\big{(}t\chi_{f}\big{)}^{2}+(3+4r)t\chi_{f}+4r^{2}+6r+e_{f}$
[…]
Hence
$\big{(}t\chi_{f}\big{)}^{2}+\big{(}(3+4r)t+12-\lambda_{f}\big{)}\chi_{f}+4r^{2}+6r\geq 2g+1>25\chi_{f}^{2}+28\chi_{f}+9.$ (4.4)
Note that $\lambda_{f}\leq 9$ by (2.5), and $t=\lambda_{f}-\frac{4(g-1)}{g}\leq\frac{5g+4}{g}$. Moreover, by (4.1) and (4.2),
$r=\frac{k_{0}(k_{0}-1)}{g}$ $\,\leq\frac{\frac{5\chi_{f}+2}{2}\cdot\frac{5\chi_{f}}{2}}{\frac{25}{2}\chi_{f}^{2}+14\chi_{f}+4}<\frac{1}{2};$
$(3+4r)t+12-\lambda_{f}$ $\,=(2+4r)\lambda_{f}-\frac{4(3+4r)(g-1)}{g}+12$
$\,\leq 9(2+4r)-\frac{4(3+4r)(g-1)}{g}+12=\frac{(16+20g)r+6g+12}{g}+12$
$\,<\frac{(16+20g)\cdot\frac{1}{2}+6g+12}{g}+12=28+\frac{20}{g}.$
Therefore, (4.4) implies that
$425\chi_{f}^{2}+28\chi_{f}+9<\bigg{(}\frac{5g+4}{g}\bigg{)}^{2}\chi_{f}^{2}+\bigg{(}28+\frac{20}{g}\bigg{)}\chi_{f}+$ 4 .
It follows that $5<\frac{40\chi_{f}^{2}+20\chi_{f}}{g}+\frac{16\chi_{f}^{2}}{g^{2}}$, which contradicts (4.1). Finally for the case when $g(B)\geq 2$, the proof would be the same as above. We simply argue as follows. Suppose that
$g>32\chi_{f}^{2}+20\chi_{f}+3.$ (4.5)
Then $s_{g+2}=0$ by (2.7). Let
$\tilde{k}_{0}=\max\left\{k~{}\big{|}~{}s_{2k}\neq 0,\text{~{}or~{}}s_{2k+1}\neq 0\right\}.$
Then by (2.7) one shows similarly that
$\tilde{k}_{0}\leq 4\chi_{f}+1.$ (4.6)
Let
$\tilde{r}=\frac{\tilde{k}_{0}(\tilde{k}_{0}-1)}{g},\quad t=\lambda_{f}-\frac{4(g-1)}{g},\quad a_{k}=\frac{4k(g-k)-g}{g},\quad b_{k}=\frac{2k(g-k+1)-2g}{g}.$
Then by (2.2),
$\big{(}t\chi_{f}\big{)}^{2}+(3+4\tilde{r})t\chi_{f}+4\tilde{r}^{2}+6\tilde{r}\geq\sum_{k=1}^{\tilde{k}_{0}}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{\tilde{k}_{0}}(4k^{2}-2k-2)s_{2k}.$
And hence
$\big{(}t\chi_{f}\big{)}^{2}+\big{(}(3+4\tilde{r})t+12-\lambda_{f}\big{)}\chi_{f}+4\tilde{r}^{2}+6\tilde{r}$
=\,= $\big{(}t\chi_{f}\big{)}^{2}+(3+4\tilde{r})t\chi_{f}+4\tilde{r}^{2}+6\tilde{r}+e_{f}$
[…]
Similarly by applying (2.3) instead of (2.5), one obtains
$464\chi_{f}^{2}+40\chi_{f}+7<\bigg{(}\frac{8g+4}{g}\bigg{)}^{2}\chi_{f}^{2}+\bigg{(}40+\frac{20}{g}\bigg{)}\chi_{f}+$ 4 .
It follows that $3<\frac{64\chi_{f}^{2}+20\chi_{f}}{g}+\frac{16\chi_{f}^{2}}{g^{2}}$, which contradicts (4.5). This completes the proof. ∎
[…]
(1) From the proof above, one can even obtain a quadratic upper bound in Theorem 4.1 (3) such that the quadratic coefficient relies on the slope $\lambda_{f}$.
(2) The quadratic upper bounds given in Theorem 4.1 (3) seem not sharp. Nevertheless, we will construct examples in the next section showing that the quadratic upper bounds can not improved into linear ones.
5. Hyperelliptic Albanese fibrations with a quadratic Albanese genus
In this section, we will construct examples showing that the genus $g$ of hyperelliptic Albanese fibrations can be as large as a quadratic function in $K_{S}^{2}$ or $\chi(\mathcal{O}_{S})$.
Theorem 5.1.
There exist a sequence of minimal irregular surfaces $S_{k}$ ($k\geq 3$ and odd) of general type with $q(S_{k})=1$, such that their Albanese maps induce a hyperelliptic fibration $f_{k}:\,S_{k}\to E$ of genus $g_{k}=\frac{(k-1)(k+3)}{2}$, and that
$\chi(\mathcal{O}_{S_{k}})=\chi_{f_{k}}=\frac{3}{2}(k-1),\qquad K_{S_{k}}^{2}=K_{f_{k}}^{2}=8k-12.$
We use notations introduced in Section 2.2. To construct the required hyperelliptic fibrations is equivalent to find appropriate data $(P,R,\delta)$ with suitable singularity indices $s_{i}$’s.
For any integer $m\geq 1$, let $G_{m}=\mathbb{Z}/{m\mathbb{Z}}$ be the cyclic group of order $m$. Assume that $\sigma\in G_{m}$ acts on $\mathbb{P}^{1}$ by $\sigma(t)=\xi t$ (where $\xi$ is any $m$-th primitive unit root), and $\sigma\in G_{m}$ acts on some elliptic curve $E_{0}$ by translation $\sigma(x)=x+p$, where $p$ is a torsion point of order $m$. Let $P:=(E_{0}\times\mathbb{P}^{1})/{G_{m}}$, where $G_{m}$ acts on $E_{0}\times\mathbb{P}^{1}$ by the diagonal action. Then we have the following commutative diagram
${E_{0}\times\mathbb{P}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Pi}$${P:=(E_{0}\times\mathbb{P}^{1})/{G_{m}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$${E_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m:1}$${E:=E_{0}/G_{m}}$
Then $h:P\rightarrow E$ is a $\mathbb{P}^{1}$-bundle over the elliptic curve $E$. In fact, $P\cong\mathbb{P}(\mathcal{O}_{E}\oplus\mathcal{N})$, and it admits two sections $C_{0},C_{\infty}$ with $C_{0}^{2}=C_{\infty}^{2}=0$, where $\mathcal{N}$ is a torsion line bundle of order $m$ over $E$. Denote by $\Gamma$ the general fiber of $h$. As a geometric ruled surface over an elliptic curve, the Picard group of $P$ is simply:
Pic ($\text{{\rm Pic\,}}(P)=\mathbb{Z}[C_{0}]\oplus h^{*}\text{{\rm Pic\,}}(E).$
[…]
Let $h:P\rightarrow E$ be the $\mathbb{P}^{1}$-bundle and above. For any irreducible curve $D\subseteq P$, let $D\sim aC_{0}+b\Gamma$. Then $a\geq 0$, $b\geq 0$, and $a+b>0$. Moreover, if $b=0$, then $a\geq m$ unless $D=C_{0}$ or $C_{\infty}$.
Proof.
We prove here that $a\geq m$ if $D\neq C_{0}$, $D\neq C_{\infty}$ and $b=0$. The rest is clear by [Har77, § V.2]. Let $D^{\prime}=\Pi^{*}(D)\subseteq E_{0}\times\mathbb{P}^{1}$. Since $D\sim aC_{0}$, it follows that $D^{\prime}\sim aC^{\prime}$, where $C^{\prime}$ stands for $E_{0}\times\{x\}\subseteq E_{0}\times\mathbb{P}^{1}$. In particular, $D^{\prime}$ consists of several sections (the number is exactly $a$) of the trivial $\mathbb{P}^{1}$-bundle $E_{0}\times\mathbb{P}^{1}\to E_{0}$. On the other hand, as $D\neq C_{0}$ and $D\neq C_{\infty}$, it follows that any irreducible component in $D^{\prime}$ maps to another component under any non-identity element of the Galois group $G_{m}$. It follows that the number of irreducible components in $D^{\prime}$ is a multiple of $m$, and in particular $a\geq m$ as required. ∎
[…]
Let $h:P\rightarrow E$ be the $\mathbb{P}^{1}$-bundle and above. For any $p\in P\setminus\{C_{0}\cup C_{\infty}\}$, let $\tau:\,\widetilde{P}\to P$ be the blowing-up centered at $p$, and $\mathcal{E}$ be the exceptional curve. Then for any odd $k\geq 3$ and $g_{k}=\frac{(k-1)(k+3)}{2}$, there exists a smooth divisor $\widetilde{R}\in\left|\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}\right|$ if $m\geq k+2$.
Proof.
It is enough to prove that the linear system $\left|\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}\right|$ is base-point-free. Note that $K_{\widetilde{P}}=\tau^{*}K_{P}+\mathcal{E}$. It follows that $\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}=K_{\widetilde{P}}+L$, where
$L=\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma-K_{P}\big{)}-(2k+1)\mathcal{E}.$
Note that numerically we have $K_{P}\sim-2C_{0}$. Hence $L^{2}=4k+3\geq 15>0$. We claim that
[…]
For any irreducible curve $\widetilde{D}\subseteq\widetilde{P}$, $L\cdot\widetilde{D}\geq 2$. In particular, $L$ is ample.
Assuming the above claim, then one sees easily that the linear system
$\left|\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}\right|=\left|K_{\widetilde{P}}+L\right|$
is base-point-free by applying Reider’s method [Rei88]. Hence a general divisor $\widetilde{R}\in\left|K_{\widetilde{P}}+L\right|$ is smooth by Bertini’s theorem. It remains to prove Section 5. Let $\widetilde{D}\subseteq\widetilde{P}$ be any irreducible curve. If $\widetilde{D}=\mathcal{E}$, or $\widetilde{D}$ is the strict transform of any fiber of $h:\,P\to E$, then one checks easily that
$L\cdot\widetilde{D}>2.$
Otherwise, $\widetilde{D}\sim\tau^{*}(aC_{0}+b\Gamma)-\beta\mathcal{E}$ with $a>0$, i.e., the restriction $h|_{\widetilde{D}}:\,\widetilde{D}\to E$ is surjective. Thus
$0\leq 2p_{a}(\widetilde{D})-2=(K_{\widetilde{P}}+\widetilde{D})\cdot\widetilde{D},\qquad\text{namely,}\qquad 2(a-1)b\geq\beta(\beta-1).$ (5.1)
By direct computation,
$L\cdot\widetilde{D}=2\big{(}a+(g_{k}+2)b\big{)}-(2k+1)\beta=2a+(k+1)^{2}b-(2k+1)\beta.$
If $\beta>k$, i.e., $\beta\geq k+1$, then
$L\cdot\widetilde{D}$ $\,=2+2(a-1)+(k+1)^{2}b-(2k+1)\beta$
$\,\geq 2+2\sqrt{2(k+1)^{2}(a-1)b}-(2k+1)\beta$
$\,\geq 2+2(k+1)\sqrt{\beta(\beta-1)}-(2k+1)\beta$
$\,>2+2(k+1)\big{(}\beta-\frac{1}{2}\big{)}-(2k+1)\beta$
$\,=1+\beta-k\geq 2;$
if $\beta\leq k$, then
$L\cdot\widetilde{D}$ $\,=2+2(a-1)+(k+1)^{2}b-(2k+1)\beta$
{$&2+2(k+1)^{2}-(2k+1)k=3k+4>2,&\quad&\text{if~{}}b\geq 2;\\ &2+\beta(\beta-1)+(k+1)^{2}-(2k+1)\beta=2+(k+1-\beta)^{2}>2,&&\text{if~{}}b=1,\\ &2,&&\text{if $b=0$ and $a=1$};\\ &2a-(2k+1)\beta\geq 2m-(2k+1)>2,&&\text{if $b=0$ and $a\geq 2$};\end{aligned}\right.$
Here we explain a little more about the case when $b=0$: if $a=1$, then $\widetilde{D}$ must be the strict transform of $C_{0}$ or $C_{\infty}$ by Section 5, which implies that $\beta=0$ since $p\in P\setminus\{C_{0}\cup C_{\infty}\}$, and hence $L\cdot\widetilde{D}=2$; if $a\geq 2$, then $\beta\leq 1$ by (5.1), and $a\geq m$ by Section 5. ∎
We can now construct the required hyperelliptic Albanese fibrations.
Proof of Theorem 5.1.
For any odd $k\geq 3$, let $\widetilde{R}\in\left|\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}\right|$ be any smooth divisor on $\widetilde{P}$ as in Section 5. Let $R=\tau(\widetilde{R})\subseteq P$ be its image in $P$. Then $R$ admits a singularity of order $2k$ at the point $p$, and
$\mathcal{O}_{P}(R)\cong\delta^{\otimes 2},\qquad\text{where~{}}\delta=\mathcal{O}_{P}\big{(}(g_{k}+1)C_{0}+\Gamma\big{)}.$
Hence one can construct a hyperelliptic fibration $f_{k}:\,S_{k}\to E$ using the above data $(P,R,\delta)$. The genus $g_{k}$ of a general fiber of $f_{k}$ is $g_{k}=\frac{(k-1)(k+3)}{2}$. Moreover, using the notations introduced in Section 2.2, one computes that $s_{2k}=1$, $s_{2}=(K_{\widetilde{P}/E}+\widetilde{R})\cdot\widetilde{R}=10k-8$, and $s_{i}=0$ for other $i$. Hence by Theorem 2.1,
$\chi(\mathcal{O}_{S_{k}})=\chi_{f_{k}}=\frac{3}{2}(k-1),\qquad K_{S_{k}}^{2}=K_{f_{k}}^{2}=8k-12.$
Suppose that $q(S_{k})=g(E)=1$. Then $f_{k}:\,S_{k}\to E$ is nothing but the Albanese fibration of $S_{k}$. Thus it remains to show that $q(S_{k})=1$. After blowing up $\tau:\,\widetilde{P}\to P$, one sees that $S_{k}$ is a double cover of $\widetilde{P}$ branched over the smooth divisor $\widetilde{R}$. Moreover, similar to Section 5, one shows that $\widetilde{R}$ is ample. Hence $q(S_{k})=q(\widetilde{P})=g(E)=1$ as required. ∎